OpenStudy (anonymous):

Find the product. -380 x 7/16 A. 280 B. 140 C. –140 D. –280

OpenStudy (flower2468):

you have my pic!!

OpenStudy (purple_pink):

i do? :/

OpenStudy (anonymous):

No, me

OpenStudy (anonymous):

Anyway.. I need an answer please ;m;

alones (alones):

Okay so to me x looks like multiplying so here how i'd solve it (−380)(7/)16 =−2660/16 =−665/4

rebeccaxhawaii (rebeccaxhawaii):

^^